OpenStudy (anonymous):

Solve using the multiplication principle -4x>1/11

OpenStudy (anonymous):

multiply by -4 and switch the inequality

OpenStudy (anonymous):

I think you mean \(-{1\over 4}\)

OpenStudy (anonymous):

nope it says -4x

OpenStudy (anonymous):

\[-4x>\frac{1}{11}\]\[x>\frac{1}{11}*\frac{-1}{4}\]\[x<\frac{-1}{44}\]

OpenStudy (anonymous):

I was talking to satellite, not you rachel ;)

OpenStudy (anonymous):

oh..lol sorry

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-4x%3E1%2F11